DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, July 10, 2003 Padelford C-36 1:30 pm (Note: Special Day and Time!)

Removable singularities: thin and fat, straight or not, old and new

Thierry De Pauw (Universite Paris Sud, France)

I will briefly review what ’removable singularities’ problems are about. The core of the talk will be about removable singularities for complex analytic functions, the Laplace equation, the minimal surface equation and the Gauss-Green Theorem. I will mention G. David and P. Mattila’s results about Lipschitz harmonic functions in the plane. Next I will explain new results obtained with W.F. Pfeffer about the minimal surface equation and the Gauss-Green Theorem, as well as a new negative result explaining why purely unrectifiable sets are not removable for the Gauss-Green formula. Finally I will try to convince you that some conjectures in the area are related to the regularity of the singular set of some class of minimal surfaces.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.